OpenStudy (gitface):

Simplify tan(sin^-1(x)).

OpenStudy (anonymous):

is that inverse?

OpenStudy (gitface):

\[\tan (\sin^{-1} (x))\]

OpenStudy (gitface):

I began by letting \[\sin^{-1} (x) = \Theta \] therefore \[\sin(\Theta) = x\] I now seek tan(theta) \[\tan(\Theta) = \frac{ \sin(\Theta) }{ \cos(\Theta) }\] and here I get stuck.

zepdrix (zepdrix):

Work from the inside out.\[\Large \arcsin x=\theta \qquad\to\qquad \sin \theta = \frac{x}{1}\]The reason I wrote it like that, is so we can relate it to the sides of a triangle. Remember the relationship that sine has with a triangle?

OpenStudy (gitface):

sinx = y/r

OpenStudy (gitface):

or opp/hyp

zepdrix (zepdrix):

Yes, but you're thinking too fancy. Ya there we go :)

zepdrix (zepdrix):

So let's set up a triangle! :)

zepdrix (zepdrix):



zepdrix (zepdrix):

Understand where we're going with this?

OpenStudy (gitface):

I think. The hyp= 1 is the radius of the unit circle, correct?

zepdrix (zepdrix):

We don't actually want to relate this back to the unit circle :) Let's find the missing side of the triangle.

OpenStudy (gitface):

You mean an expression for the bottom side? \[\sqrt{x ^{2}-1}\]

zepdrix (zepdrix):

I think you have that backwards maybe :o